OpenStudy (anonymous):

find the angle 0 which determines the given length.Express results in degrees, accurate to two decimal places. r=10 ft,s=30 ft How would I solve this.

OpenStudy (benjieg6pd):

O is the measure of a central angle subtended by arc s=30 ft on a circle with radius r=10ft. First find the measure of angle O in radians by using the formula s=(r)(O) : 30 = 10 * O Divide both sides by 10 O = 3 radians Convert 3 radians to degrees: O = 3 * (180/pi) Pi = 3.14 approx O = 3 * (180/3.14) O = 171.97 degrees